OpenStudy (anonymous):

What would be the resulting molarity of a solution made by dissolving 31.3 grams of Ca(OH)2 in enough water to make a 1050-milliliter solution?

OpenStudy (australopithecus):

\[Molarity = \frac{Moles}{Litres}\] 1L = 1000mL \[Moles = \frac{Grams}{Molecular Mass}\] Note Molecular mass has units g/mol

OpenStudy (australopithecus):

You can find the molecular mass of a compound by looking at your periodic table and adding the weights up for each element in the compound for example the molar mass of HCl is Molar Mass of H + Molar Mass of Cl 1.00794g/mol + 35.453g/mol = Molar Mass of Cl

OpenStudy (australopithecus):

Molar Mass = Molecular Mass Any questions ask or if you get stuck show me your work and I can help you

OpenStudy (australopithecus):

I pretty much just gave you the answer though, these questions are plug and chug questions you should memorize both those formulas btw

OpenStudy (anonymous):

Ok

OpenStudy (australopithecus):

Just to avoid confusion if you want to find the molecular mass of H2O for example it would be Molar mass of H + Molar Mass of H + Molar Mass of O = Molar Mass of H2O

OpenStudy (anonymous):

So it would be 40.078/mol + (15.999*1.008)2/mol... like this

OpenStudy (anonymous):

(15.999*1.008)+(15.999*1.008)

OpenStudy (australopithecus):

the units are g/mol not 1/mol secondly you need to add the molecular mass of each atom so in your case Ca(OH)2 broken down Ca + OH + OH you have the atoms Ca + H + O + O + H add their molecular mass

OpenStudy (australopithecus):

to get the molecular mass of the compound

OpenStudy (australopithecus):

its like if you had 12 apples all had different weights it doesn't make any sense to multiply their weights, you just need to add them all up.

OpenStudy (australopithecus):

Molar mass is essentially the weight of an atom per 1 mole of it. so 1 mole of calcium = 40.078grams

OpenStudy (australopithecus):

we want the weight of the entire compound per 1 mole, to do that we have to add them all up

OpenStudy (australopithecus):

a mole being avagadros number of atoms or molecules

OpenStudy (anonymous):

ok so 40.078 + 1.008 + 15.998 +1.008 +15.998

OpenStudy (anonymous):

Which is 79.04

OpenStudy (anonymous):

i meant 74.09

OpenStudy (australopithecus):

looks good to me, make sure to use the units g/mol

OpenStudy (australopithecus):

showing units is important

OpenStudy (anonymous):

Then do we divide this

OpenStudy (anonymous):

ok do we put something over this or am I just over thinking it

OpenStudy (australopithecus):

nvm no you are right lol

OpenStudy (australopithecus):

keep going convert to moles then, convert mL to Liters and plug into molarity formula

OpenStudy (australopithecus):

\[Moles = \frac{Grams of Ca(OH)_2}{79.04g/mol}\]

OpenStudy (anonymous):

74.09/100=0.7409